THE INVERSION SPLITTING OF ^{15}NH_{2}D AND ^{15}ND_{2}H AS OBTAINED FROM THEIR FIR SPECTRA

M. ELKEURTI, Laboratoire d’Etudes Physico Chimiques, 20000 Saïda, Algérie; \underline{L. H. COUDERT}, J. ORPHAL, LISA, CNRS/Universités Paris 12 et 7, 61 Avenue du Général de Gaulle, 94010 Créteil, France; C. E. FELLOWS, Laboratório de Espectroscopia e Laser, Universidade Federal Fluminense, Campus da Boa Viagem, Niterói, RJ, 24210-340, Brazil; AND S. TOUMI, Institut d’Electronique, Université Mokhtar Badji, 23000 Annaba, Algérie.

The far infrared absorption spectra of the partially deuterated species ^{15}NH_{2}D and ^{15}ND_{2}H have been recorded in the 20 to 105 cm^{−1} region using a Brucker IFS 120 Fourier transform spectrometer. A 25 cm long absorption cell with teflon windows was filled with different mixtures of ^{15}NH_{3} and ^{15}ND_{3} at total pressures ranging from 0.8 to 2 mbar. Several spectra with different ^{15}NH_{2}D to ^{15}ND_{2}H ratio were thus recorded. For all spectra the maximum path length was used resulting in a 2 × 10^{−3} cm^{−1} resolution. Prior to the line assignment, rotational constants for ^{15}NH_{2}D and ^{15}ND_{2}H where calculated from those of ^{14}NH_{2}D and 14ND_{2}H. The IAM-type theoretical approach developed by Cohen and Pickett,^{a} which accounts for the large amplitude inversion motion and its coupling with the overall rotation, was afterwards used to obtain rovibrational energies and predicted spectra. In the case of ^{15}NH_{2}D, observed line positions for low-J lines were sometimes within 0.050 cm^{−1} from their predicted values. The line position analyzes are still in progress. For both species, they are carried out with the help of the IAM-type theoretical approach.^{a} For ^{15}NH_{2}D more than 300 a- and c-type lines have already been assigned and the observed wavenumbers are reproduced with an RMS deviation of 0.4 × 10^{−3} cm^{−1}. This will be discussed in the paper and the values obtained for the spectroscopic constants will be reported and compared to those obtained by Fusina et al.^{b} for the 14N-species. The decrease of the tunneling inversion splitting which goes from 0.406 cm^{−1} in ^{14}NH_{2}D^{b} to 0.383 cm^{−1} in 15NH_{2}D will also be discussed.

aE. A. Cohen and H. M. Pickett, J. Molec. Spectrosc. 93, 83 (1982).

bL. Fusina, G. Di Lonardo, J. W. C. Johns and L. Halonen, J. Molec. Spectrosc. 127, 240 (1988).